OpenStudy (zasharra):

I really need help! It won't take long. You must: Label and display your new polynomial identity Prove that it is true through an algebraic proof, identifying each step Demonstrate that your polynomial identity works on numerical relationships (x – y) (x + y) (y + x) (y – x) (x + a) (y + b) (x2 + 2xy + y2) (x2 – 2xy + y2) (ax + b) (cy + d)

OpenStudy (zasharra):

@welshfella do you know where to begin?

OpenStudy (welshfella):

I guess it wants you to pick out an identity from some of the above expressions one would be (x + y)^2 = x^2 + 2xy + y^2 if you expand the left side using the distributive law you will get the right side

OpenStudy (welshfella):

thats the proof . Then you can plug in some values of x and y to demonstrate the last part.

OpenStudy (zasharra):

@welshfella dude you totally stole that.. smh you could have gotten me in trouble!!! I REALLY DO NEED HELP. if i change those 2's into 5's do you know how carry on with the problem??

OpenStudy (zasharra):

(x + y)^5 = x^5 + 5xy + y^5

OpenStudy (zasharra):

@welshfella

OpenStudy (welshfella):

no that is not correct

OpenStudy (zasharra):

i really dont know how to to do this. can you give an original example

OpenStudy (welshfella):

the right side is much longet YOu need to know the Binomial Theorem to get that one.

OpenStudy (zasharra):

???

OpenStudy (welshfella):

OK Prove the identity ( a - 2b)^2 = a^2 - 4ab + b^2 applying the distributive law LHS = (a - 2b)(a - 2b) = a(a - 2b) - 2b(a - 2b) = a^2 - 2ab - 2ab + b^2 = a^2 - 4ab + b^2 = RHS that is the algebraic proof

OpenStudy (welshfella):

now plug in values like a = 2 and b = 3 and find if LHS = RHS for 3rd part

OpenStudy (welshfella):

LHS = left hand side RHS = right hand side

OpenStudy (welshfella):

oh inade one error the last term is 4b^2 not b^2

OpenStudy (zasharra):

so what would be the final answer

OpenStudy (welshfella):

for the proof where i have b^2 replace it with 4b^2 you can do the last part just put a=2 and b = 3 into (a - 2b)^2 = a^2 - 4ab + 4b^2 to show that left side = right side left side = (2 -2*3)^2 = -4^2 = 16 so see if you can get right side = 16 It should do.

OpenStudy (zasharra):

ok thank you so much

OpenStudy (welshfella):

yw